xt Problem 9 Suppose that the economys production function is given by where Y is output, K is capital, and N is the number of workers. The steady-state level of capital per worker in terms of the saving rate, s, and the depreciation rate, 8, is ON=□ Property format your oxpression using the too sinthe palette Hover over tools t see keyboard shortcuts. E character) as person t can be Enter your arawer in the answer box and then click Check Answer parts Ciedr All Check Answer MacBook Air 3 4 5 7 K.
Answer #1

Y/N = \sqrt{K}\sqrt{N}

Y/N = \sqrt{K/N}\sqrt{N/N}

Y/N = \sqrt{K/N}.

The following equation shows capital per worker from year t to year (t+1)

Kt+1/N - Kt/N = (s) f(K*/N) - (\delta) (K*/N)

Under steady state, the capital stock does not change, so Kt+1/N - Kt/N = 0.

(s) f(K*/N) - (\delta) (K*/N) = 0

(s) f(K*/N) = (\delta) (K*/N)

s * \sqrt{K/N} = \delta (K/N)

Squaring both sides,

s2 (K/N) = \delta2 (K/N)2

s2/\delta2 = (K/N).